# FedMeld: A Model-dispersal Federated Learning Framework for Space-ground Integrated Networks

### v1

@@ II-C Performance Metric
At the $t$-th iteration, the convergence rate of FedMeld algorithm is defined as the expectation of the global loss function at the $t$-th step, i.e., $F\left({{{\overline{\mathbf{w}}}_{t}}}\right)$, and the optimal objective value $F\left({{{\mathbf{w}}^{\star}}}\right)$, as
@@ III FedMeld Algorithm and Convergence Analysis
In this section, we first describe our proposed FedMeld algorithm. Then, we conduct convergence analysis of FedMeld under full and partial participation schemes. The results are useful for optimizing the FedMeld in the next section.
@@ III-A FedMeld Algorithm
We divide the serving satellites into SCF satellites and non-SCF satellites. The former is the serving satellites which carry the latest models from regions until reaching the next serving area. The latter also acts as parameter servers, performs FedAvg for regions before the SCF satellites come, and does not store and carry models. For $t\in\mathcal{I}_{E}$, there are two cases:
@@
2) When the serving satellite is an SCF satellite, area $(i-1)$’s model carried by this satellite can be mixed with area $i$’s model after performing FedAvg for the collected clients’ model in step $t$.
Without loss of generality, we assume the handover occurs only at synchronization steps. To accelerate the training process, devices in each region continue additional training rounds with non-SCF satellites before the arrival of the incoming SCF satellite. We define the global round interval of model mixing between adjacent regions as $\delta$. During the stage from when the SCF satellite leaves the visible range of area $i$ until the completion of model aggregation for area $\left(i+1\right)$, we define two parameters: $T_{i,i+1}^{{\rm{fly}}}$ and $T_{i,i+1}^{{\rm{idle}}}$. Here, $T_{i,i+1}^{{\rm{fly}}}$ is the total duration of this process, and $T_{i,i+1}^{{\rm{idle}}}$ is the idle duration of clients in the region $(i+1)$, i.e., they are neither training with SCF satellites nor non-SCF satellites. Then, we have $T_{i,i+1}^{{\rm{idle}}}+\sum\limits_{k^{\prime}=k+1}^{k+\delta}{T_{k^{\prime},i+1}^{{\rm{total}}}}=T_{i,i+1}^{{\rm{fly}}}$ for any $k-K\bmod(K+\delta)=0$. Since we consider that all the regions are covered by a specific circular orbit, area $i=0$ is equivalent to the last area (i.e., $i=M$), and $i=M+1$ equals the first area (i.e., $i=1$).
Let a binary variable, $A_{t}\in\left\{0,1\right\}$, denote whether the area’s model mixes with its adjacent model at step $t$, which is predetermined by the network topology of SGINs. Let $\alpha$ denote the model mixing ratio of adjacent areas. Then, the model of device $j\in\mathcal{N}_{i}$ at step $t$ can be expressed as
${\mathbf{w}_{t,j}}=\left\{\begin{array}[]{l}{\mathbf{v}_{t,j}},\;{\text{if}}\;t\notin{\mathcal{I}}_{E},\\ \left({1-\alpha{A_{t}}}\right){{\overline{\mathbf{v}}}_{t,i}}+{\alpha}{A_{t}}{{\overline{\mathbf{w}}}_{t-{\delta E},i-1}},\;{\text{if}}\;t\in{\mathcal{I}}_{E}.\end{array}\right.$ (9)
Fig. 2 provides an example of how models from four regions are mixed. Here, red, blue, yellow, and green stars are SCF satellites that carry the latest models from four areas, i.e., A, B, C, and D, respectively, and will perform model mixing upon reaching the next serving area. Gray satellites are non-SCF satellites and can also act as parameter servers when an area is not covered by an SCF satellite. Their participation leads to model staleness when mixing. Algorithm 1 shows the procedure of FedMeld, and the details are as follows.
@@
    At the beginning of a mix cycle, users in each area and their serving (nearest) SCF satellite execute $K$ global rounds cooperatively until the handover happens. Specifically, when a global round starts, the SCF satellite selects a set of participating clients to conduct $E$ local iterations. Then, the satellite collects the uploaded models from these users, runs FedAvg, and broadcasts the averaged model back to users.
@@
    When the SCF satellite serves for area $i$, users in region $i$ conduct $E$ local iterations and upload their models to this satellite. After running FedAvg for the collected models of area $i$, the satellite mixes this averaged result with its stored historical model of area $\left(i-1\right)$, where the mixing proportion of the historical model is $\alpha$.
@@
Figure 2: An example diagram of parameter mixing in FedMeld algorithm.
In the FedMeld algorithm, each time an SCF satellite arrives at the next region, a parameter mixing occurs with adjacent regional models. Since these regions are covered by satellites on the same orbital plane, only $M-1$ mixing times are required for each region to incorporate parameter information from other regions, naturally achieving global model fusion with FedMeld.
@@
III-B Assumptions and Definitions
We proceed by making the widely adopted assumptions on client $j$’s local loss function ${F_{j}}\left({\bf{w}}\right)$. Let $\nabla{F_{j}}\left({\bf{w}}\right)\buildrel\Delta\over{=}{{\mathbb{E}}_{{\varsigma_{j}}}}\left[{\nabla{F_{j}}\left({{\bf{w}};{\varsigma_{j}}}\right)}\right]$.
Assumption 1.
(L-smoothness): For device $j\in{\mathcal{N}}$, ${F_{j}}\left({\bf{w}}\right)$ is differentiable and there exists a constant $L\geq 0$ such that for all $\bf{w}$ and $\bf{w}^{\prime}$, ${F_{j}}\left(\mathbf{w}\right)\leq{F_{j}}\left(\mathbf{w}^{\prime}\right)+{\left({\mathbf{w}-\mathbf{w^{\prime}}}\right)^{T}}\nabla{F_{j}}\left(\mathbf{w^{\prime}}\right)+\frac{L}{2}{\left\|{\mathbf{w}-\mathbf{w^{\prime}}}\right\|^{2}}$ holds.
Assumption 2.
($\mu$-strongly convex): For device $j\in{\mathcal{N}}$, ${F_{j}}\left(\mathbf{w}\right)\geq{F_{j}}\left(\mathbf{w}^{\prime}\right)+{\left({\mathbf{w}-\mathbf{w^{\prime}}}\right)^{T}}\nabla{F_{j}}\left(\mathbf{w^{\prime}}\right)+\frac{\mu}{2}{\left\|{\mathbf{w}-\mathbf{w^{\prime}}}\right\|^{2}}$ holds.
Assumption 3.
(Unbiasedness and bounded stochastic gradient variance): For device $j\in{\mathcal{N}}$, the stochastic gradient is unbiased with variance bounded by $\sigma^{2}$ such that ${\mathbb{E}}\left[\nabla{F_{j}}\left({{\mathbf{w}}_{j};{\varsigma_{j}}}\right)\right]=\nabla{F_{j}}\left({\mathbf{w}}_{j}\right)$ and $\mathbb{E}_{\varsigma_{j}}\left\|\nabla F_{j}\left({\mathbf{w}}_{j};\varsigma_{j}\right)-\nabla F_{j}({\mathbf{w}}_{j})\right\|^{2}\leq\sigma_{j}^{2}$ for all ${\mathbf{w}}$.
Assumption 4.
(Bounded gradient): The expected value of the squared norm of stochastic gradients is consistently bounded: ${{{\left\|\nabla{F_{j}}\left({{\mathbf{w}}_{j};{\varsigma_{j}}}\right)\right\|}^{2}}}\leq G^{2}$ for all $j\in\mathcal{N}$.
@@ III-B Assumptions and Definitions
Definition 1.
(Degree of non-IID) Let $F^{*}$ and $F_{j}^{*}$ be the minimum values of $F$ and $F_{j}$, respectively. We use a constant $\Gamma$, denoted by $\Gamma={F^{*}}-\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}\frac{1}{N_{i}}{\sum\limits_{j\in{\mathcal{N}}_{i}}{F_{j}^{*}}}$, to quantify the non-IID degree of data in SGINs.
If the data are IID, $\Gamma$ goes to zero when the number of data increases. If the data are non-iid, we will have $\Gamma>0$ and its magnitude shows the data heterogeneity.
III-C Convergence Analysis
We first analyze the convergence bound of FedMeld under the full participation scheme and then extend the result to the partial participation one.
@@ III-C Convergence Analysis
(Results of one step SGD [26, Lemma 1]) Let assumption 1 and 2 hold. Notice that ${{\overline{\mathbf{v}}}_{t+1}}={{\overline{\mathbf{w}}}_{t}}-\eta_{t}{\mathbf{g}_{t}}$ always holds. If $\eta_{t}\leq\frac{1}{4L}$, we have
@@
Define function $m\left(\alpha\right)={\frac{{4\left({1-\alpha}\right)\left({2{\alpha^{2}}-\alpha+1}\right)}}{{\alpha\left({2{\alpha^{2}}-3\alpha+3}\right)}}}$. With Lemma 1, we have the following theorem to provide the convergence bound of FedMeld under full participation scheme.
Theorem 1.
(Convergence bound under full participation scheme) Choose the learning rate $\eta_{t}=\frac{5}{\mu(\gamma+t)}$ for $\gamma=\max\left\{{\frac{{8L}}{\mu},E}\right\}-1$, when $\delta E\leq\frac{{m\left(\alpha\right){{\left({t+\gamma+1}\right)}^{2}}}}{{{{\left({t+\gamma}\right)}^{2}}+m\left(\alpha\right)\left({t+\gamma+1}\right)}}$, the convergence bound with full participation at the last time step $R$ satisfies
$\begin{split}&\mathbb{E}\left[{F\left({{{\overline{\mathbf{w}}}_{R}}}\right)}\right]-F\left({{{\mathbf{w}}^{\star}}}\right)\leq\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)\frac{{L}}{{\mu\left({R+\gamma}\right)}}\\ &\cdot\left[{\frac{{25B}}{{{8\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}\mathbb{E}{{\left\|{{{\overline{\mathbf{w}}}_{1}}-{{\mathbf{w}}^{\star}}}\right\|}^{2}}}\right]+L{Q_{R-1}},\end{split}$ (11)
where ${Q_{R-1}}=\mathbb{E}\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{{\mathcal{N}}_{i}}}{\frac{1}{{M{N_{i}}}}}}{\left\|{{\mathbf{w}_{{R-1},j}}-{{\overline{\mathbf{w}}}_{R-1}}}\right\|^{2}}$ and $B=\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{{\mathcal{N}}_{i}}}{\frac{{\sigma_{j}^{2}}}{{{{\left({M{N_{i}}}\right)}^{2}}}}}}+6L\Gamma$.
@@
For any $t\geq 0$ satisfying $t\in\mathcal{I}_{E}$, $Q_{t}=\mathbb{E}\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{\mathcal{N}_{i}}}{\frac{1}{{M{N_{i}}}}}}{\left\|{{\mathbf{w}_{t,j}}-{{\overline{\mathbf{w}}}_{t}}}\right\|^{2}}$ has the following upper bound,
@@
Based on Lemma 2, the term $Q_{R-1}$ can be upper bounded as follows.
@@
Assume that ${R-1}$ is the final step of a particular mix cycle. The upper bound of $Q_{R-1}$ satisfies
@@
After completing the analysis of the full participation scheme, we extend this result to the partial participation scenario as follows.
@@
(Bounding average variance under unbiased sampling scheme [29, Lemma 1]) If $t$ is a global round and satellites select clients randomly without replacement, then $\mathbb{E}_{\mathcal{U}_{t}}\overline{\mathbf{z}}_{t}=\overline{\mathbf{w}}_{t}$, where $\mathcal{U}_{t}=\bigcup_{i\in\mathcal{M}}\mathcal{U}_{t,i}$ represents the union of selected clients. Due to partial participation, we have
@@
Based on Lemma 3, the convergence bound under the partial participation scheme is provided in the following theorem.
Theorem 3.
(Convergence Bound under Partial Participation) With an unbiased sampling scheme, the convergence bound with partial client participation at step $R$ can be expressed as
@@
Remark 1.
(Impact of partial participation): The partial participation scheme will bring an additional term reflected by $\mathop{\max}\limits_{i\in\mathcal{M}}{\frac{{{N_{i}}-{U_{i}}}}{{M{U_{i}}\left({{N_{i}}-1}\right)}}}$. When $U_{i}$ is larger, the negative effect caused by this term will be mitigated. In the extreme case that all clients join the training process, this term will be eliminated.
Remark 2.
(Discussion of $b$ in Theorem 2): It is clear that $\kappa_{1}$ has the following upper bound, i.e., ${\kappa_{1}}\leq{\left({1+b}\right)^{K+\delta}}{c^{2}}\cdot\max\left\{{{\alpha^{2}},{{\left({1-\alpha}\right)}^{2}}}\right\}$, where $c\geq 1$. Define an auxiliary variable $\rho$ satisfying $1<\rho\leq\min\left\{{\frac{1}{\alpha},\frac{1}{{1-\alpha}}}\right\}$. When choosing $b={\rho^{\frac{1}{{K+1}}}}-1$ and $c=\frac{1}{{\rho\cdot\max\left\{{\alpha,1-\alpha}\right\}}}\geq 1$ for $\delta<K+2$, $\kappa_{1}<1$ can be guaranteed. Then, $\kappa_{1}$ and $\kappa_{2}$ can be recast as
[…]
@@
Remark 3.
(Impact of $\delta$ and $\alpha$ on model accuracy): Given the communication scenario and initial training setting in SGINs, $E$, $G$, and $K$ (since the training rounds of each serving satellite are stable in the long term) are fixed. Therefore, the model loss in (16) can be adjusted by designing parameters $\delta$ and $\alpha$. We can find that the right side of (16) is a monotonically increasing function of $\delta$, indicating the tradeoff between latency and accuracy. However, the coupling relationship between $\delta$ and $\alpha$ not only exists in the upper bound of model loss but also in the constraint presented in (1). Their complex dependency makes it challenging to derive their joint optimal solutions, motivating us to formulate an optimization problem and design an algorithm in the next section.

### v2

@@ II-C Performance Metric
At the $t$-th iteration, the convergence rate of FedMeld algorithm is defined as the difference between the expectation of the global loss function at the $t$-th step, i.e., $F\left({{{\overline{\mathbf{w}}}_{t}}}\right)$, and the optimal objective value $F\left({{{\mathbf{w}}^{\star}}}\right)$, as
@@
For the general partial participation scheme, the global virtual sequence $\overline{\mathbf{w}}_{t}$ is substituted by ${\overline{\mathbf{z}}_{t}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{\frac{1}{{{U_{i}}}}\sum\limits_{j\in{{\mathcal{U}}_{t,i}}}{{{\mathbf{w}}_{t,j}}}}$.
@@ III FedMeld Algorithm and Convergence Analysis
In this section, we first describe our proposed FedMeld algorithm. Then, we conduct a convergence analysis of FedMeld under full and partial participation schemes. The results will be used for optimizing the FedMeld in the next section.
@@ III-A FedMeld Algorithm
To realize infrastructure-free model aggregation, FedMeld classifies serving satellites into two functional roles based on their capability to carry and forward models across regions:
  • SCF satellites: They are responsible for inter-regional model mixing by storing aggregated models from one region and carrying them to the next along their orbital path. Upon arrival at a new region, they mix the stored model from the previous region with the newly aggregated model from local clients, enabling infrastructure-free parameter mixing across adjacent regions.
  • Non-SCF satellites [^2] The spatial distribution of ground clusters may be uneven, leading to variations in the flight interval between adjacent regions. The number of global training rounds $K$ for each region is jointly determined by the satellite constellation configuration, the geographical size of each region, the spatial distance between adjacent regions, and the user–satellite communication latency. When two regions are geographically close, the time left for the satellite to serve the next region after completing the current one may be limited. Therefore, $K$ should be carefully designed considering the above factors to ensure that the satellite can complete $K$ rounds within each coverage period. Conversely, when regions are far apart, the idle interval can be effectively utilized by non-SCF satellites, which continue local training before receiving the model from the previous region.: They act as temporary parameter servers that perform local model aggregation (FedAvg) for the current region only. These satellites do not store or forward models across regions and simply discard the aggregated models after completing the local service. As a result, their operation introduces asynchronous model updates across regions, since clients in one area can continue local training while awaiting the arrival of an SCF satellite carrying models from previous regions.
Remark 1.
(Necessity of Non-SCF Satellites) The primary role of non-SCF satellites is to utilize the idle time when an SCF satellite flies from one region to the next. Without non-SCF satellites, clients in a region must wait until the SCF satellite arrives to perform model mixing, resulting in synchronous training and idle periods. In contrast, the presence of non-SCF satellites enables asynchronous model mixing, allowing clients in the next region to continue training and aggregating models before an SCF satellite carrying the model from the previous region arrives. It is worth noting that designating all satellites as SCF is not desirable. In such a case, multiple outdated models would be repeatedly forwarded and mixed across regions even when fresher models are already available, which diminishes their contribution to convergence. This SCF-based division aligns with the orbital movement of satellites, where only selected satellites, determined by their orbital paths and the order in which they serve regions, are designated to carry forward models between consecutive regions.
For $t\in\mathcal{I}_{E}$, there are two cases:
@@
2) When the serving satellite is an SCF satellite, area (i1)(i-1)’s model carried by this satellite will be mixed with area $i$’s model after performing FedAvg for the collected clients’ model in step $t$.
Without loss of generality, we assume that handover occurs after the previous serving satellite broadcasts the latest averaged model to the clients. To accelerate the training process, devices in each region continue additional training rounds with non-SCF satellites before the arrival of the incoming SCF satellite. We define the global round interval of model mixing between adjacent regions as $\delta$, indicating that when model mixing occurs at step $t$, area $i$’s model at step $t$ will be mixed with area (i1)(i-1)’s model from step $t-\delta E$. Since we consider that all the regions are covered by a specific circular orbit, area $i=0$ is equivalent to the last area (i.e., $i=M$), and $i=M+1$ equals the first area (i.e., $i=1$).
Fig. 2 provides an example of how models from four regions are mixed. The procedure of FedMeld is presented below.
@@
    At the beginning of a mix cycle, users in each area and their serving (nearest) SCF satellite execute $K$ global rounds cooperatively until the handover happens. Within each global round, the SCF satellite selects a set of participating clients to conduct $E$ local iterations, collects the uploaded models from these users, runs FedAvg, and broadcasts the averaged model back to users. The client selection adopts a random without-replacement strategy, ensuring that each client participates at most once per round and that the global update remains unbiased in expectation.
@@
    When the SCF satellite serves area $i$, users in region $i$ conduct $E$ local iterations and upload their models to this satellite. After running FedAvg for the collected models of area $i$, the satellite mixes this averaged result with its stored historical model of area $\left(i-1\right)$, where the mixing proportion of the historical model is $\alpha$. Let a binary variable, $A_{t}\in\left\{0,1\right\}$, denote whether the area’s model mixes with its adjacent model at step $t$. When $t\mod(K+\delta)=0$, $A_{t}=1$ holds, otherwise, $A_{t}=0$. Therefore, the model of device $j\in\mathcal{N}_{i}$ can be updated as follows:
[…]
@@
Figure 2: A diagram of parameter mixing in FedMeld algorithm. Colored satellites represent SCF satellites that carry aggregated models across adjacent regions and perform inter-region mixing upon arrival. Gray satellites denote non-SCF satellites, which only perform local aggregation (FedAvg) during their service period without model transfer to subsequent regions.
The FedMeld algorithm is summarized in Algorithm 1. In FedMeld, each time an SCF satellite arrives at the next region, a parameter mixing occurs with the adjacent regional models. As a result, after M1M-1 mixing, each region incorporates parameter information from all other regions, naturally achieving global model fusion with FedMeld.
@@
III-B Convergence Analysis
Intuitively, FedMeld results in model parameter exchanges across regions by satellite mobility. However, does it converge, and how quickly does it converge? This requires a convergence analysis of FedMeld.
We proceed by making the widely adopted assumptions on client $j$’s local loss function ${F_{j}}\left({\bf{w}}\right)$. Let $\nabla{F_{j}}\left({\bf{w}}\right)\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}{{\mathbb{E}}_{{\varsigma_{j}}}}\left[{\nabla{F_{j}}\left({{\bf{w}};{\varsigma_{j}}}\right)}\right]$.
Assumption 1 (L-smoothness).
For device $j\in{\mathcal{N}}$, ${F_{j}}\left({\bf{w}}\right)$ is differentiable and there exists a constant $L\geq 0$ such that for all $\bf{w}$ and $\bf{w}^{\prime}$, ${F_{j}}\left(\mathbf{w}\right)\leq{F_{j}}\left(\mathbf{w}^{\prime}\right)+{\left({\mathbf{w}-\mathbf{w^{\prime}}}\right)^{T}}\nabla{F_{j}}\left(\mathbf{w^{\prime}}\right)+\frac{L}{2}{\left\|{\mathbf{w}-\mathbf{w^{\prime}}}\right\|^{2}}$ holds.
Assumption 2 ($\mu$-strongly convex).
For device $j\in{\mathcal{N}}$, ${F_{j}}\left(\mathbf{w}\right)\geq{F_{j}}\left(\mathbf{w}^{\prime}\right)+{\left({\mathbf{w}-\mathbf{w^{\prime}}}\right)^{T}}\nabla{F_{j}}\left(\mathbf{w^{\prime}}\right)+\frac{\mu}{2}{\left\|{\mathbf{w}-\mathbf{w^{\prime}}}\right\|^{2}}$ holds.
Assumption 3 (Unbiasedness and bounded stochastic gradient variance).
For device $j\in{\mathcal{N}}$, the stochastic gradient is unbiased with variance bounded by $\sigma^{2}$ such that ${\mathbb{E}}\left[\nabla{F_{j}}\left({{\mathbf{w}}_{j};{\varsigma_{j}}}\right)\right]=\nabla{F_{j}}\left({\mathbf{w}}_{j}\right)$ and $\mathbb{E}_{\varsigma_{j}}\left\|\nabla F_{j}\left({\mathbf{w}}_{j};\varsigma_{j}\right)-\nabla F_{j}({\mathbf{w}}_{j})\right\|^{2}\leq\sigma_{j}^{2}$ for all $\mathbf{w}$}.
Assumption 4 (Bounded gradient).
The expected value of the squared norm of stochastic gradients is consistently bounded: ${{{\left\|\nabla{F_{j}}\left({{\mathbf{w}}_{j};{\varsigma_{j}}}\right)\right\|}^{2}}}\leq G^{2}$ for all $j\in\mathcal{N}$.
@@ III-B Convergence Analysis
Definition 1 (Degree of non-IID).
Let $F^{*}$ and $F_{j}^{*}$ be the minimum values of $F$ and $F_{j}$, respectively. We use a constant $\Gamma$, denoted by $\Gamma={F^{*}}-\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}\frac{1}{N_{i}}{\sum\limits_{j\in{\mathcal{N}}_{i}}{F_{j}^{*}}}$, to quantify the non-IID degree of data in SGINs. If the data are IID, $\Gamma$ goes to zero when the number of data increases. If the data are non-iid, we will have $\Gamma>0$ and its magnitude shows the data heterogeneity.
Based on the assumptions above, we first analyze the convergence bound of FedMeld under the full participation scheme and then extend the result to the partial participation one. For analysis, we introduce an auxiliary variable $\mathbf{v}_{t+1,j}$} to denote the intermediate result of one step stochastic gradient descent (SGD) update from $\mathbf{w}_{t,j}$}. That is,
${\mathbf{v}_{t+1,j}}={\mathbf{w}_{t,j}}-{\eta_{t}}{\nabla{F_{j}}\left({\mathbf{w}_{t,j},\varsigma_{t,j}}\right)}.$ (9)
Then, the update rule of the model parameter of client $j$ under FedMeld can be rewritten as follows:
${\mathbf{w}_{t,j}}=\begin{cases}{\mathbf{v}_{t,j}}&\text{ if }t\notin\mathcal{I}_{E}\\ \left({1-\alpha{A_{t}}}\right){{\overline{\mathbf{v}}}_{t,i}}+{\alpha}{A_{t}}{{\overline{\mathbf{w}}}_{t-{\delta E},i-1}},&\text{ if }t\in\mathcal{I}_{E}\end{cases}.$ (10)
We also define a virtual sequence as ${{\overline{\mathbf{v}}}_{t}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{{{\overline{\mathbf{v}}}_{t,i}}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{\frac{1}{{{N_{i}}}}\sum\limits_{j\in{{\mathcal{N}}_{i}}}{{{\mathbf{v}}_{t,j}}}}$, where ${\overline{\mathbf{v}}}_{t,i}$}} is area $i$’s virtual sequence.
To facilitate the subsequent convergence proof, we first recall a classical result of one-step SGD, which serves as the foundation for our later derivations.
@@
(Results of one step SGD [42, Lemma 1]) Let Assumption 1 and 2 hold. Notice that ${{\overline{\mathbf{v}}}_{t+1}}={{\overline{\mathbf{w}}}_{t}}-\eta_{t}{\mathbf{g}_{t}}$ always holds. If $\eta_{t}\leq\frac{1}{4L}$, we have
@@
With Lemma 1, we have the following theorem to provide the convergence bound of FedMeld under full participation scheme.
Theorem 1 (Convergence bound under full participation scheme).
Choose the learning rate $\eta_{t}=\frac{5}{\mu(\gamma+t)}$ for $\gamma=\max\left\{{\frac{{8L}}{\mu},E}\right\}-1$, when $\delta E\leq\frac{{m\left(\alpha\right){{\left({t+\gamma+1}\right)}^{2}}}}{{{{\left({t+\gamma}\right)}^{2}}+m\left(\alpha\right)\left({t+\gamma+1}\right)}}$, the convergence bound with full participation at the last time step $R$ satisfies
$\begin{split}\mathbb{E}\left[{F\left({{{\overline{\mathbf{w}}}_{R}}}\right)}\right]&-F\left({{{\mathbf{w}}^{\star}}}\right)\leq\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)\frac{{L}}{{\mu\left({R+\gamma}\right)}}\\ &\cdot\left[{\frac{{25B}}{{{8\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}\mathbb{E}{{\left\|{{{\overline{\mathbf{w}}}_{1}}-{{\mathbf{w}}^{\star}}}\right\|}^{2}}}\right]+L{Q_{R-1}},\end{split}$ (12)
where $m\left(\alpha\right)={\frac{{4\left({1-\alpha}\right)\left({2{\alpha^{2}}-\alpha+1}\right)}}{{\alpha\left({2{\alpha^{2}}-3\alpha+3}\right)}}}$, ${Q_{R-1}}=\mathbb{E}\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{{\mathcal{N}}_{i}}}{\frac{1}{{M{N_{i}}}}}}{\left\|{{\mathbf{w}_{{R-1},j}}-{{\overline{\mathbf{w}}}_{R-1}}}\right\|^{2}}$ and $B=\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{{\mathcal{N}}_{i}}}{\frac{{\sigma_{j}^{2}}}{{{{\left({M{N_{i}}}\right)}^{2}}}}}}+6L\Gamma$.
@@
The constraint $\delta E\leq\frac{{m\left(\alpha\right){{\left({t+\gamma+1}\right)}^{2}}}}{{{{\left({t+\gamma}\right)}^{2}}+m\left(\alpha\right)\left({t+\gamma+1}\right)}}$ in Theorem 1 imposes an upper bound on the round interval $\delta$, which quantifies the staleness of models exchanged between adjacent regions. Practically, this ensures that the model carried by the SCF satellite is not excessively outdated when mixed with the latest model in the next region. If $\delta$ exceeds this limit, the stale models will lead to performance degradation and hinder convergence.
To analyze the evolution of model divergence across clients and clusters, we define $Q_{t}=\mathbb{E}\sum\limits_{i\in{\mathcal{M}}}{\sum\limits_{j\in{\mathcal{N}_{i}}}{\frac{1}{{M{N_{i}}}}}}{\left\|{{\mathbf{w}_{t,j}}-{{\overline{\mathbf{w}}}_{t}}}\right\|^{2}}$, which measures the expected variance of local models with respect to the global mean at step $t$. This quantity serves as a key indicator in the subsequent analysis. With this definition, we next establish an upper bound for $Q_{t}$, which forms the basis for the overall convergence proof.
@@
For any $t\geq 0$ satisfying $t\in\mathcal{I}_{E}$, $Q_{t}$ has the following upper bound,
@@
Based on the recursive relationship in Lemma 2, we can now derive an upper bound of $Q_{R-1}$, as summarized in Theorem 2.
@@
Assume that R1{R-1} is the final step of a particular mix cycle. The upper bound of $Q_{R-1}$ satisfies
@@
Having established the convergence bound under full participation, we extend our analysis to the more practical case of partial client participation.
@@
(Bounding average variance under unbiased sampling scheme [43, Lemma 1]) If $t$ is a global round and satellites select clients randomly without replacement, then $\mathbb{E}_{\mathcal{U}_{t}}\overline{\mathbf{z}}_{t}=\overline{\mathbf{w}}_{t}$, where $\mathcal{U}_{t}=\bigcup_{i\in\mathcal{M}}\mathcal{U}_{t,i}$ represents the union of selected clients. Due to partial participation, we have
@@
Using the above Lemma 3, the convergence bound under the partial participation scheme can be derived as follows.
Theorem 3 (Convergence bound under partial participation).
With an unbiased sampling scheme, the convergence bound with partial client participation at step $R$ can be expressed as
$\mathbb{E}\left[{F\left({{{\overline{\mathbf{z}}}_{R}}}\right)}\right]-F\left({{{\mathbf{w}}^{\star}}}\right)\leq{\zeta_{1}}\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)+{\zeta_{2}}\frac{{{\kappa_{2}}}}{{1-{\kappa_{1}}}}+{\zeta_{3}}.$
@@
Building upon the convergence result established in Theorem 3, we next provide several remarks to interpret its implications and practical insights.
Remark 2 (Impact of partial participation).
The partial participation scheme will bring an additional term reflected by $\mathop{\max}\limits_{i\in\mathcal{M}}{\frac{{{N_{i}}-{U_{i}}}}{{M{U_{i}}\left({{N_{i}}-1}\right)}}}$. When $U_{i}$ is larger, the negative effect caused by this term will be mitigated. If all clients join the training process, this term will be eliminated.
Remark 3 (Discussion of $b$ in Theorem 2).
It is clear that $\kappa_{1}$ has the following upper bound, i.e., ${\kappa_{1}}\leq{\left({1+b}\right)^{K+\delta}}{c^{2}}\cdot\max\left\{{{\alpha^{2}},{{\left({1-\alpha}\right)}^{2}}}\right\}$, where $c\geq 1$. Define an auxiliary variable $\rho$ satisfying $1<\rho\leq\min\left\{{\frac{1}{\alpha},\frac{1}{{1-\alpha}}}\right\}$. When choosing $b={\rho^{\frac{1}{{K+1}}}}-1$ and $c=\frac{1}{{\rho\cdot\max\left\{{\alpha,1-\alpha}\right\}}}\geq 1$ for $\delta<K+2$, $\kappa_{1}<1$ can be guaranteed. Then, $\kappa_{1}$ and $\kappa_{2}$ can be recast as
$\left\{\begin{aligned} \kappa_{1}&=\rho^{\frac{K+\delta}{K+1}}(1-\alpha)^{2}+\rho\alpha^{2},\\ \kappa_{2}&=4{\left({E-1}\right)^{2}}{G^{2}}\frac{{{\rho^{\frac{1}{{K+1}}}}}}{{{\rho^{\frac{1}{{K+1}}}}-1}}\left[{\frac{{{\rho^{\frac{{K+\delta}}{{K+1}}}}{{\left({1-\alpha}\right)}^{2}}+\rho{\alpha^{2}}}}{{{\rho^{\frac{1}{{K+1}}}}-1}}+{{\left({1-\alpha}\right)}^{2}}}\right].\end{aligned}\right$.
@@
Remark 4 (Impact of $\delta$ and $\alpha$ on model accuracy).
Given the communication scenario and initial training setting in SGINs, $E$, $G$, and $K$ (since the training rounds of each serving satellite are stable in the long term) are fixed. Therefore, the upper bound of the training loss in (17) can be adjusted by designing parameters $\delta$ and $\alpha$. We can find that the right side of (17) is a monotonically increasing function of $\delta$, indicating the tradeoff between latency and accuracy. However, the coupling relationship between $\delta$ and $\alpha$ not only exists in the upper bound of the training loss but also in the constraint presented in (1). Their complex dependency makes it challenging to derive their joint optimal solutions, motivating us to formulate an optimization problem and design an algorithm in the next section.